If the price of an apple increases from P20.00 to P25.00 and the amount you buy fails from 20 to 16 apple then your electricity of demand would be?
1
Expert's answer
2021-10-06T09:45:14-0400

"Price elasticity of demand =\frac{ percentage change in Quantity}{percentage change in price}"


"Percentage change in price = 25-20= 5"


"= \frac{5}{20}×100= 25"%

"Percentage change in Quantity = 16-20= -4" "= \frac{-4}{16}×100= -25"



"Price elasticity of demand = \frac{-25}{25}""= -1"